OpenStudy (anonymous):

simplify 5p/3q-4p^2/q

OpenStudy (whpalmer4):

Make a common denominator, subtract, then simplify.

OpenStudy (anonymous):

with which

OpenStudy (whpalmer4):

\[\frac{5p}{3q} - \frac{4p^2}{q}\] Why don't you multiply the right hand fraction by \[\frac{3}{3}\] and see where that gets you?

OpenStudy (anonymous):

k